OpenStudy (anonymous):

What is 13/20 written as a decimal

OpenStudy (anonymous):

0.52

OpenStudy (anonymous):

Thanks....!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

which one... A. 0.55 B. 0.65 C. 0.60 D. 0.13

OpenStudy (anonymous):

Oh sorry I miscalculated, it should be B, 0.65.

OpenStudy (anonymous):

Thank you and Its ok....Again thanks lots..

OpenStudy (radar):

13/20 = 6.5/10 =.65

OpenStudy (anonymous):

Thank you alllll.....zzzz!!

OpenStudy (anonymous):

A bag contains 14 golf balls and 16 tennis balls. What is the ratio, written in lowest terms, of golf balls to the total number of balls in the bag? A. 7 to 8 B. 7 to 15 C. 8 to 15 D. 7 to 16 Question Resources

OpenStudy (anonymous):

Can you help me on this last one as well....??

OpenStudy (radar):

First you need the total number of balls.-----Addition Then you set up a fraction the numerator equal to the number of golf balls, and the denominator equal to the total number of balls. Reduce this fraction, then write this "numerator to denominator" putting in the appropriate numbers. Then examine your choices and choose the one that matches.